(a)   A survey of 25 individuals who passed the ACCA exam showed that average annual salary is RM150,000 and standard deviation RM15,000.


(i)                    Find the degrees of freedom.

 

(ii)                  Construct 99% confidence interval for the true average salary.



1
Expert's answer
2021-08-05T13:56:50-0400

(i) "df=n-1=25-1=24" degrees of freedom.

(ii) The critical value for "\alpha=0.01" and "df=24" degrees of freedom is "t_c=2.79694."

The corresponding confidence interval is computed as shown below:


"CI=(x-t_c\times\dfrac{s}{\sqrt{n}},x+t_c\times\dfrac{s}{\sqrt{n}} )"

"=(150000-2.79694\times\dfrac{15000}{\sqrt{25}},"

"150000+2.79694\times\dfrac{15000}{\sqrt{25}})"

"=(141609.18, 158390.82)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "141609.18<\mu<158390.82," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(141609.18, 158390.82)."